A canoeist rows downstream in 1.5 hours and back upstream in 3 hours. What is the rate of the canoe if the distance rowed is 12 miles in each direction?

One-fifth of a hive of bees flew to the Kadamba flower; one-third flew to the Silandhara; three times the difference of these two numbers flew to an arbor, and one bee continued flying about, attracted on each side by the fragrant Keteki and the Malati. Find the number of bees.

Answer by mananth(16946)

Speed in still water = x mph
speed of current = y mph
against current x-y 3.00 hours
with current x+y 1.50 hours

Distance = same= 12
t=d/r
12 / ( x - y )= 3.00
3 ( x - y ) = 12.00
3 x -3 y = 12 ....................1
12 / ( x + y )= 1.50
1.50 ( x + y ) = 12
1.50 x + 1.50 y = 12 ...............2
Multiply (1) by 1
Multiply (2) by 2
we get
3 x + -3 y = 12
3 x + 3 y = 24
6 x = 36
/ 6
x = 6 mph Boat speed

Let the number of bees be x
1/5 x Kadamba
1/3 x Silandhara
3(1/3 x -1/5 x ) arbor
1 bee no flower
x/5 +x/3 +3(x/3-x/5)+1 =x
x/5+x/3+x-3x/5 +1 =x
LCD = 15
multiply by 15
3x+5x+15x-9x+15=15x
-x+15=15
-x=-15
x=15 bees